# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v5

@@ Abstract
Li-CO2 batteries show promise as energy storage solutions, offering high theoretical energy density and CO2 fixation. Their operation is based on the formation and decomposition of Li2CO3/C during discharge and charge cycles, respectively. We used a multiscale modeling framework that integrates Density Functional Theory (DFT), Ab-Initio Molecular Dynamics (AIMD), classical Molecular Dynamics (MD), and Finite Element Analysis (FEA) to investigate atomic and cell-level properties. The considered Li-CO2 battery consists of a lithium metal anode, an ionic liquid electrolyte, and a carbon cloth porous cathode with Sb0.67Bi1.33Te3 as a catalyst. DFT and AIMD determined the electrical conductivities of Sb0.67Bi1.33Te3 and Li2CO3 using the Kubo-Greenwood formalism and studied the CO2 reduction mechanism on the cathode catalyst. MD simulations calculated the CO2 diffusion coefficient, Li+ transference number, ionic conductivity, and Li+ solvation structure. The FEA model, incorporating results from atomistic simulations, reproduced experimental voltage-capacity profiles at 1 mA/cm2 and revealed spatio-temporal variations in Li2CO3/C deposition, porosity, and CO2 concentration dependence on discharge rates in the cathode. Accordingly, Li2CO3 can form large and thin film deposits, leading to dispersed and local porosity changes at 0.1 mA/cm2 and 1 mA/cm2, respectively. The capacity decreases exponentially from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, due to pore clogging from excessive discharge product deposition that limits CO2 transport to the cathode interior. Therefore, the performance of Li-CO2 batteries can be improved by enhancing CO2 transport, regulating Li2CO3 deposition, and optimizing cathode architecture.
@@ 1 Introduction
Lithium-carbon dioxide (Li-CO2) batteries, with their high theoretical energy density and CO2 fixation capability, have emerged as a potential energy storage solution [1, 2]. These batteries provide a theoretical specific energy density of 1876 Wh kg-1, which is substantially higher than that of current lithium-ion battery systems [3, 4]. This high energy density makes Li-CO2 batteries particularly attractive for applications in aviation, aerospace, and electric vehicles, where lighter and more compact energy storage solutions are crucial [5]. The potential for electric vehicles with much longer ranges or considerably lighter batteries could revolutionize the automotive market [6, 5]. However, Li-CO2 batteries are still in early development stages and face challenges including reaction reversibility issues, low energy efficiency, and the need for improved cathode designs [7]. Ongoing research and development efforts focus on overcoming these limitations, with recent advancements showing promising results [4, 8, 9, 10, 11, 12, 13, 14]. The Li-CO2 batteries operate on a unique electrochemical principle involving the reversible reaction between lithium and carbon dioxide [1, 2] contributing to CO2 capture and utilization efforts [11, 15]. During the discharge process, lithium oxidizes at the anode, releasing lithium ions and electrons . Besides, at the cathode in the presence of solid catalysts [16, 4], CO2 undergoes reduction (), combining with lithium ions and electrons to form primarily lithium carbonate (Li2CO3) and carbon (C) . This discharge reaction typically occurs at a voltage of approximately 2.8 V [17]. The charging process reverses these reactions, decomposing the discharge products to regenerate lithium and CO2. However, due to the stability of Li2CO3, this process requires higher voltages, often exceeding 3.8 V [18]. This voltage requirement highlights one of several key challenges facing Li-CO2 battery technology. The formation of solid Li2CO3 and C can lead to electrode clogging and increased internal resistance [9, 19]. The high charging voltages necessary for Li2CO3 decomposition reduce overall energy efficiency [20, 21]. Effective cathode catalysts are critical for enhancing reaction kinetics and minimizing overpotentials in Li-CO2 batteries [8, 22]. Unwanted side reactions, particularly electrolyte decomposition, can occur at high voltages [7]. Furthermore, the inherent inertness of CO2 makes its electrochemical reduction challenging, affecting discharge voltage and efficiency [23, 2, 15].
The Li-CO2 battery chemistry faces critical challenges in catalyst performance, primarily characterized by two fundamental issues, catalyst inefficiency and redox mediator limitations. Noble and transition metal catalysts (such as Pt and Au nanoparticles) demonstrate low catalytic activities for CO2 reduction/evolution reactions in aprotic media, while 2D materials (e.g., MoS2) exhibit poor structural stability at high current rates [18]. Existing catalysts struggle to maintain consistent performance across multiple charge-discharge cycles, often due to catalyst deactivation caused by self-degradation or environmental factors such as discharge product accumulation and species contamination [18]. Unlike Li-O2 batteries, Li-CO2 battery liquid catalysts are typically inactive in decomposing discharge products (Li2CO3 and solid carbon), and current redox mediators fail to effectively facilitate the reversible transformation of reaction intermediates [24]. The formation of Li2CO3, a wide-bandgap insulator, leads to slow kinetics and high voltage platforms (> 4.3 V) during the charging process, significantly decreasing the round-trip efficiency of the battery [25]. The lack of efficient decomposition mechanisms significantly impedes battery cycling efficiency, with most current Li-CO2 batteries only capable of operating for dozens of cycles with low overpotential, even with state-of-the-art catalysts [18]. These challenges necessitate developing novel catalytic systems that can maintain high catalytic activity across extended cycling, demonstrate structural stability at elevated current rates, enable efficient decomposition of discharge products, and provide reversible electrochemical reactions with minimal energy loss. The development of catalysts with high catalytic activity and conductivity is crucial to the research of high-performance Li-CO2 batteries [22]. Promising research directions in Li-CO2 battery technology encompass a range of innovative approaches aimed at overcoming current limitations and enhancing overall performance [26, 25, 27]. Advanced solid redox mediators have shown significant potential, as demonstrated by a study using a Cu(II) coordination compound of benzene-1,3,5-tricarboxylic acid. This solid redox mediator enabled a Li-CO2 battery to achieve a higher discharge voltage of 2.8 V, a lower charge potential of 3.7 V, and superior cycling performance over 400 cycles [26]. The use of solid redox mediators addresses the shuttle effect and sluggish kinetics associated with soluble redox mediators, potentially improving long-term stability and efficiency. Alternative reaction pathway engineering has emerged as another promising direction, focusing on promoting the formation of Li2C2O4 instead of Li2CO3 as the discharge product. This approach can reduce the charge potential and improve overall battery performance [25]. For instance, the use of Mo2C as a cathode material has been shown to stabilize Li2C2O4 as the final discharge product, reducing the charge potential to 3.8 V and enhancing the reversibility of the battery [25]. Innovative catalyst designs, including carbon-based and metal-complex catalysts, are being extensively explored. Carbon-based materials, such as multi-walled carbon nanotubes (MWCNTs), are widely used as cathodes due to their high electrical conductivity and large surface area [27]. However, their limited catalytic activity for Li2CO3 decomposition has led to investigations of ways to enhance their performance, such as introducing metal catalysts like Ru to reduce overpotentials and improve efficiency [27]. Transition metal compound catalysts have shown promise in improving the catalytic activity and stability of Li-CO2 batteries. For example, the integration of Ru-catalyzed MWCNT cathodes with solid electrolytes has demonstrated enhanced electrochemical reactions involving CO2 [27]. Additionally, nanostructured catalyst development with improved conductivity and catalytic properties continues to be a focus area, exploring various materials and structures to optimize battery performance [4, 10, 11].
Ionic liquid-based electrolytes show promise for Li-CO2 batteries due to their stability with intermediate products compared to solid-state electrolytes, but their development is limited by low ionic conductivity and high viscosity [18]. To address these limitations, ionic liquid electrolytes based on DMSO (dimethyl sulfoxide) as a solvent and LiTFSI (lithium bis(trifluoromethanesulfonyl)imide) as a salt, often incorporating EMIM-BF4 (1-Ethyl-3-methylimidazolium tetrafluoroborate), have been developed [28, 29, 30, 31, 3, 32, 4]. These electrolytes are preferred for Li-CO2 batteries due to their superior ionic conductivity, enhanced electrochemical stability, and improved CO2 activation properties. DMSO-based electrolytes exhibit faster Li+ ion diffusion kinetics compared to other solvents like TEGDME (tetraethylene glycol dimethyl ether), resulting in lower polarization potential [28]. Li-CO2 batteries using 3M LiTFSI in DMSO electrolyte have demonstrated excellent cycling performance, maintaining stable operation for over 1,900 hours (180 cycles) [28]. In concentrated DMSO-based electrolytes, TFSI- anions can access the primary solvation sheath of Li+ ions, leading to the formation of a LiF-rich solid electrolyte interphase (SEI) layer, which is beneficial for battery performance [29, 30]. DMSO-based electrolytes demonstrate good electrochemical stability with both the Li metal anode and the reactive cathode at high voltages and elevated temperatures [28]. The use of DMSO-based electrolytes improves the rate capability of the battery, with cells able to deliver flat charge/discharge plateaus at current densities of up to 2,000 mA g-1 [29]. Together, these properties make the DMSO LiTFSI electrolyte a promising choice for improving the overall performance and stability of Li-CO2 batteries. The addition of EMIM-BF4 (1-Ethyl-3-methylimidazolium tetrafluoroborate) as an electrolyte material plays several important roles in Li-CO2 batteries. EMIM-BF4 helps in activating CO2 molecules, which is crucial for the CO2 reduction reaction in Li-CO2 batteries [31]. The addition of EMIM-BF4 to the electrolyte has been shown to enhance battery performance. In one study, a Li-CO2 battery using 0.1 M LiTFSI in (EMIM-BF4)/DMSO as the electrolyte achieved 500 cycles at 0.5 A/g with a capacity of 500 mAh/g [28]. The presence of EMIM-BF4 in the electrolyte contributes to improved cycling stability of Li-CO2 batteries [3]. When combined with other electrolyte components like DMSO and LiTFSI, EMIM-BF4 contributes to the overall performance enhancement of Li-CO2 batteries, including higher discharge voltages and increased capacities [32]. Recently, Jaradat et al. used an electrolyte combination of ZnI2, LiTFSI, DMSO, and EMIM-BF4 with a volumetric ratio of 2:3 that led to good Li-CO2 battery performance [4].
Although recent research has focused on novel cathode materials and electrolytes to improve Li-CO2 battery cell performance [10, 33, 34, 11, 14, 13, 35], most computational work has focused primarily on atomistic-level or continuum-level modeling. There is a critical need for a comprehensive multiscale approach that bridges these disparate levels of analysis to provide a more holistic understanding of batteries [36, 37, 38, 39]. To address this gap, we present an interactive multiscale modeling framework for Li-CO2 batteries that integrates density functional theory (DFT), ab initio molecular dynamics (AIMD), classical molecular dynamics (MD), and finite element analysis (FEA) in a coordinated manner. This approach bridges the atomic properties and electrochemical performance of Li-CO2 batteries.
@@
At the Li-CO2 battery cell level, Finite Element Analysis (FEA) effectively simulates charge/discharge cycles, enabling macroscopic analysis of complex processes within the battery cell. These processes include ion transport through porous electrodes, reaction kinetics, and their impact on cell performance [40, 41, 42]. However, FEA requires accurate electrode and electrolyte properties as input parameters, which can be derived from atomistic simulations. At the atomic scale, classical Molecular Dynamics (MD) simulations excel in estimating transport properties such as diffusion coefficients and ionic conductivity in battery electrolytes [43, 44]. Complementing this approach, the combination of Density Functional Theory (DFT) and Ab Initio Molecular Dynamics (AIMD) enables the calculation of electrode electrical conductivity and reaction rate coefficients at the electrolyte/electrode interface [45, 12, 46, 47].
To bridge the gap between atomic and cell-level simulations, we first employ DFT/AIMD calculations to determine the electrical conductivity of Li2CO3 and Sb0.67Bi1.33Te3 using the Kubo-Greenwood formula [48, 49, 50], as well as to study the surface reaction of Sb0.67Bi1.33Te3 as cathode catalyst material. Secondly, we conducted large-scale MD simulations to investigate the transport properties of the liquid electrolyte. Specifically, we calculate the diffusion coefficients of Li+ and CO2, and the Li+ transference number by examining the mean square displacement (MSD) of the particles over time using the Einstein equation [51], and the ionic conductivity using the Green-Kubo formula [52]. Finally, the atomic-scale insights gained from DFT/AIMD and MD simulations are then integrated into our FEA model, allowing for a comprehensive analysis of cell-level performance. Our FEA simulations investigate various aspects of battery cell behavior, including specific capacity, discharge behavior, and voltage-capacity profiles considering factors such as CO2 transport properties, cathode porosity, and deposition of Li2CO3 and C. The studied Li-CO2 battery cell consists of a lithium anode, an ionic liquid electrolyte, and a porous cathode. The electrolyte consists of a mixture of 1-ethyl-3-methylimidazolium tetrafluoroborate (EMIM-BF4) and dimethyl sulfoxide (DMSO) in a 2:3 ratio, supplemented with zinc iodide (ZnI2) and 1 M lithium bis(trifluoromethanesulfonyl)imide (LiTFSI) salt. This specific electrolyte composition has been experimentally demonstrated to yield optimal battery cell performance [4]. The cathode uses carbon cloth coated with a stable transition metal trichalcogenide alloy catalyst (Sb0.67Bi1.33Te3), improving CO2 reduction and oxidation kinetics [4]. A schematic representation of our multiscale approach is depicted in Figure 1. This multiscale approach enables us to make predictions about porosity, deposited species quantities, and CO2 concentration variations-parameters that are difficult to assess through experimental methods alone. Furthermore, it allows us to evaluate these properties as a function of time, providing dynamic insights into the battery’s behavior. This computational methodology can be used to test and screen various battery materials, potentially saving significant time and resources by reducing the need for extensive experimental tests.
@@ 2 Methods and computational details
2.1 Density Function Theory calculations (DFT) for structure optimization
We performed first-principles-based Density Functional Theory (DFT) calculations to optimize the crystal structures of Sb0.67Bi1.33Te3 and Li2CO3 compounds using the Vienna Abinitio Simulation Package (VASP) [53, 54]. The Sb0.67Bi1.33Te3 structure is constructed based on the Sb2Te3 structure from the Materials Project database [55]. Multiple possible structures for Sb0.67Bi1.33Te3 were optimized, and the most energetically favorable configuration (i.e., with the lowest optimization energy) is presented in Figure 1 (B). The Li2CO3 structure presented in 1B is obtained from the Materials Project database [55]. The projector augmented wave (PAW) formalism described the valence electrons of Bi, Sb, Te, Li, C, and O atoms using plane wave-based wave functions was employed [56]. The structure optimization, involving the minimization of ground state energy, utilized the generalized gradient approximation of Perdew and Wang method that brings the correct exchange-correlation part of electronic energy [57, 58]. A kinetic cutoff energy of 500 eV was set to enhance calculation accuracy. The unit cell of both Sb0.67Bi1.33Te3 and Li2CO3 compounds (Figure 1 (B)) extended in all three directions such that the lattice constants become more than 20Å to sample k-meshes of the Brillouin zones, respectively. Ionic and electronic optimizations were alternately performed until the forces on each ion reached less than meV/Å.
For the Sb0.67Bi1.33Te3 surface interaction (Figure 1 (D)), the Perdew-Burke-Ernzerhof (PBE) functional, within the generalized gradient approximation (GGA), was employed to describe the exchange-correlation energy. A plane-wave basis set was used with an appropriate energy cutoff to ensure energy convergence, and the interaction between valence and core electrons was treated using the projector augmented wave (PAW) method. A cut-off energy was set at 500 eV and the Brillouin zone sampled at gamma point with a k-mesh of 4 x 4 x 1. The convergence criterion of the total energy was set to be within 1x10-5 eV within the K-point integration and optimizations were performed until the forces on each atom were below 0.001 eV/Å. Finally, to account for the van der Waals interactions, the van der Waals D3 corrections were implemented [59]. Zero-point energy, heat capacity, and entropy corrections were computed using statical mechanics within the harmonic approximation.
@@ 2.2 Ab initio molecular dynamics (AIMD) and Kubo-Greenwood method for electrical conductivity calculation
Using optimized structures of Sb0.67Bi1.33Te3 and Li2CO3 compounds from our DFT calculation (Figure 1 (B)), we carried out the Ab initio molecular dynamics (AIMD) simulations at 300K up to 1ps with NVT ensemble and 1fs time step to select 10 different frames in the 1ps trajectory data. Subsequently, we generated wavefunctions ( where and are band and k-point indices, respectively) data for the selected frames using DFT calculations as explained by Knyazev et. al. [60, 48] to compute the electronic conductivity based on Kubo-Greenwood (KG) formula. During the calculation of wavefunction data from VASP, we included 100% more unoccupied bands than occupied bands for both Sb0.67Bi1.33Te3 and Li2CO3 compounds to account for excited state wavefunctions and their corresponding energies. The dynamic electrical conductivity is represented as a coefficient of the current density and electric field :

### v6

@@ Abstract
Li-CO2 batteries are promising energy storage systems due to their high theoretical energy density and CO2 fixation capability, relying on reversible Li2CO3/C formation during discharge/charge cycles. We present a multiscale modeling framework integrating Density Functional Theory (DFT), Ab-Initio Molecular Dynamics (AIMD), classical Molecular Dynamics (MD), and Finite Element Analysis (FEA) to investigate atomic and cell-level properties. The considered Li-CO2 battery consists of a lithium metal anode, an ionic liquid electrolyte, and a carbon cloth cathode with Sb0.67Bi1.33Te3 catalyst. DFT and AIMD determined the electrical conductivities of Sb0.67Bi1.33Te3 and Li2CO3 using the Kubo-Greenwood formalism and studied the CO2 reduction mechanism on the cathode catalyst. MD simulations calculated the CO2 diffusion coefficient, Li+ transference number, ionic conductivity, and Li+ solvation structure. The FEA model, parameterized with atomistic simulations data, reproduced the available experimental voltage-capacity profile at 1 mA/cm2 and revealed spatio-temporal variations in Li2CO3/C deposition, porosity, and CO2 concentration dependence on discharge rates in the cathode. Accordingly, Li2CO3 can form large and thin film deposits, leading to dispersed and local porosity changes at 0.1 mA/cm2 and 1 mA/cm2, respectively. The capacity decreases exponentially from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, due to pore clogging from excessive discharge product deposition that limits CO2 transport to the cathode interior. Therefore, the performance of Li-CO2 batteries can be improved by enhancing CO2 transport, regulating Li2CO3 deposition, and optimizing cathode architecture.
@@ 1 Introduction
Lithium-carbon dioxide (Li-CO2) batteries, with their high theoretical energy density and CO2 fixation capability, have emerged as a potential energy storage solution [1, 2]. These batteries provide a theoretical specific energy density of 1876 Wh kg-1, which is substantially higher than that of current lithium-ion battery systems [3, 4]. This high energy density makes Li-CO2 batteries particularly attractive for applications in aviation, aerospace, and electric vehicles, where lighter and more compact energy storage solutions are crucial [5]. The potential for electric vehicles with much longer ranges or considerably lighter batteries could revolutionize the automotive market [6, 5]. However, Li-CO2 batteries are still in early development stages and face challenges including reaction reversibility issues, low energy efficiency, and the need for improved cathode designs [7]. Ongoing research and development efforts focus on overcoming these limitations, with recent advancements showing promising results [4, 8, 9, 10, 11, 12, 13, 14]. The Li-CO2 batteries operate on a unique electrochemical principle involving the reversible reaction between lithium and carbon dioxide [1, 2] contributing to CO2 capture and utilization efforts [11, 15]. During the discharge process, lithium oxidizes at the anode, releasing lithium ions and electrons . Besides, at the cathode in the presence of solid catalysts [16, 4], CO2 undergoes reduction (), combining with lithium ions and electrons to form primarily lithium carbonate (Li2CO3) and carbon (C) . This discharge reaction typically occurs at a voltage of approximately 2.8 V [17]. The charging process reverses these reactions, decomposing the discharge products to regenerate lithium and CO2. However, due to the stability of Li2CO3, this process requires higher voltages, often exceeding 3.8 V [18]. This increased voltage requirement underscores one of several key challenges facing Li-CO2 battery technology. In particular, the formation of solid Li2CO3 and carbon can lead to electrode clogging and increased internal resistance [9, 19]. Moreover, the high charging voltages needed to decompose Li2CO3 significantly reduce the overall energy efficiency of the system [20, 21]. To address these issues, effective cathode catalysts are critical for enhancing reaction kinetics and minimizing overpotentials in Li-CO2 batteries [8, 22]. Nevertheless, unwanted side reactions, particularly electrolyte decomposition, can occur at these high voltages [7]. Furthermore, the inherent inertness of CO2 compounds the challenge, as it makes electrochemical reduction difficult and adversely affects both discharge voltage and energy efficiency [23, 2, 15].
In this work, we consider a Li-CO2 battery cell comprising a lithium anode, a hybrid ionic liquid electrolyte, and a porous cathode. The electrolyte features a unique composition of 1-ethyl-3-methylimidazolium tetrafluoroborate (EMIM-BF4) and dimethyl sulfoxide (DMSO) in a 2:3 ratio, supplemented with zinc iodide (ZnI2) and 1 M lithium bis(trifluoromethanesulfonyl)imide (LiTFSI) salt. This optimized electrolyte formulation, experimentally validated for superior performance [4], enables efficient ion transport and stability. The cathode utilizes carbon cloth coated with a novel transition metal trichalcogenide alloy catalyst (Sb0.67Bi1.33Te3), which significantly enhances CO2 reduction/oxidation kinetics and cycling stability [4]. While recent advancements in Li-CO2 batteries have focused on developing novel cathode materials and electrolytes [10, 24, 25, 11, 14, 13, 26], most computational studies have been limited to atomistic-level investigations. To address this gap, we introduce the first comprehensive multiscale framework that integrates quantum-scale simulations, Molecular Dynamics, and continuum modeling of Li-CO2 batteries. This approach uniquely connects atomic-scale phenomena, such as catalyst surface reactions and ion transport, to macroscopic cell performance, enabling holistic optimization of Li-CO2 battery systems. This strategy is also needed for other battery types [27, 28, 29, 30]. Our methodology advances the field by addressing critical knowledge gaps in the spatiotemporal evolution of parameters and the interdependencies between electrochemical processes at different scales, and it can be extended to the study of other battery systems.
@@
At the Li-CO2 battery cell level, Finite Element Analysis (FEA) can effectively simulate charge/discharge cycles, enabling macroscopic analysis of complex processes within the battery cell. These processes include ion transport through porous electrodes, reaction kinetics, and their impact on cell performance [31, 32, 33]. However, FEA requires accurate electrode and electrolyte properties as input parameters, which can be derived from atomistic simulations.
To bridge the gap between atomic and cell-level simulations and to calculate the necessary parameters for FEA, we employed a closed-loop workflow from the atomic to the cell level, as summarized in Figure 1. We employ quantum mechanics (QM) and Density Functional Theory (DFT) calculations to determine partial charges, equilibrated geometries, and Molecular Dynamics Force Field (FF) parameters for battery constituents using small systems (number of atoms < 1000). Ab Initio Molecular Dynamics (AIMD) calculations are used to determine the electrical conductivity of Li2CO3 and Sb0.67Bi1.33Te3, based on DFT-equilibrated geometries and employing the Kubo-Greenwood formula [34, 35, 36]. AIMD is also used to study the surface reactions of Sb0.67Bi1.33Te3 as a cathode catalyst material. Large-scale MD simulations are conducted to investigate the transport properties of the liquid electrolyte. Specifically, we calculate the diffusion coefficients of Li+ and CO2, as well as the Li+ transference number, by examining the mean square displacement (MSD) of the particles over time using the Einstein equation [37], and the ionic conductivity using the Green-Kubo formula [38]. The atomic-scale insights gained from DFT/AIMD and MD simulations are then integrated into our FEA model, enabling a comprehensive analysis of cell-level performance. Our FEA simulations investigate various aspects of battery cell behavior, including specific capacity, discharge behavior, and voltage-capacity profiles, considering factors such as CO2 transport properties, cathode porosity, and the deposition of Li2CO3 and C.
The Li-CO2 continuum model (CM) of the Li-CO2 battery that we developed, supported by accurately calculated parameters from atomistic-level simulations, enables us to make predictions about porosity, deposited species quantities, and CO2 concentration variations—parameters that are difficult to assess through experimental methods alone. Furthermore, it allows us to evaluate these properties as a function of time, providing dynamic insights into the battery’s behavior at the cell level. This hierarchical, cross-validated workflow not only bridges the longstanding gap between atomistic understanding and macroscopic battery behavior but also enables targeted material screening and iterative optimization of cell components. The resulting predictive capability accelerates the discovery and engineering of Li-CO2 batteries with enhanced energy density, cycle life, and operational safety. Moreover, our methodology provides a generalizable template for the rational design of other advanced battery chemistries, where multiscale interactions govern performance and reliability. Ultimately, this multiscale computational strategy offers a powerful tool to guide experimental efforts, reduce development time, and unlock the full potential of emerging energy storage technologies.
@@ 2 Methods and computational details
2.1 Density Functional Theory Calculations for Structure Optimization
First-principles Density Functional Theory (DFT) calculations were performed using the Vienna Ab-initio Simulation Package (VASP) [39, 40] to optimize the crystal structures of Sb0.67Bi1.33Te3 and Li2CO3. The Sb0.67Bi1.33Te3 structure was constructed based on Sb2Te3 from the Materials Project database [41], and multiple configurations were optimized to identify the most energetically favorable structure. The Li2CO3 structure was also obtained from the Materials Project database. For surface interaction studies, the Perdew-Burke-Ernzerhof (PBE) functional within the generalized gradient approximation (GGA) was used, and dispersion interactions were accounted for using the D3 van der Waals correction [42]. These calculations are crucial for accurately modeling CO2 adsorption and reduction processes on the catalyst surface. Further computational details are provided in the Supplementary Information.
@@ 2.2 Ab Initio Molecular Dynamics (AIMD) and Kubo-Greenwood method for electrical conductivity calculation
Using optimized structures of Sb0.67Bi1.33Te3 and Li2CO3 compounds from our DFT calculation, we carried out the Ab Initio Molecular Dynamics (AIMD) simulations at 300K up to 1ps with NVT ensemble and 1fs time step to select 10 different frames in the 1ps trajectory data. Subsequently, we generated wavefunctions ( where and are band and k-point indices, respectively) data for the selected frames using DFT calculations as explained by Knyazev et. al. [43, 34] to compute the electronic conductivity based on Kubo-Greenwood (KG) formula. During the calculation of wavefunction data from VASP, we included 100% more unoccupied bands than occupied bands for both Sb0.67Bi1.33Te3 and Li2CO3 compounds to account for excited state wavefunctions and their corresponding energies. The dynamic electrical conductivity is represented as a coefficient of the current density and electric field :